I don't really know how to type this problem but I need help with this inequality | 4t-3 | is less than or equal to 7
Answer by stanbon(75887)

| 4t-3 | is less than or equal to 7
----
|4t-3| = 4t-3 if t >= 3/4
|4t-3| = 3-4t if t < 3/4
-------------------------------
Solve if t >= 3/4
4t-3 <= 7
4t <= 10
t <= 10/4
--
Solutions where (3/4)<= t <=(10/4)
========================================
Solve if t < 3/4
3-4t <= 7
4t >= -4
t >=-1
---
Combining those you get -1 <= t <= 10/4
====================================
Cheers,
Stan H.
============